OpenStudy (anonymous):

Math Help! Given the following system of equations, identify the type of system. y = -x + 4 2x + 2y = 8 independent inconsistent equivalent

OpenStudy (nikato):

hint: multiply both sides by2 for the first equation. then move the x term to the left side. what do you get?

OpenStudy (studygurl14):

Or you can divide the second equation by 2 and move the x 2x + 2y = 8 x + y = 4 y = -x + 4 Hey, that looks awfully familiar... You get these two equations y = -x + 4 y = -x + 4 They are identical, so teh answer is C

OpenStudy (anonymous):

oh, so do you divide by two for all equations

OpenStudy (nikato):

no. you could either multiply by 2 on both sides of the first equaiton. or divide by 2 on both sides for the second equation

OpenStudy (studygurl14):

Way #1: y = -x + 4 --> 2y = -2x + 8 --> 2x + 2y + 8 2x + 2y = 8 The two are identical Way #2: y = -x + 4 2x + 2y = 8 --> x + y = 4 --> y = -x + 4 The two are identical